OpenStudy (anonymous):

can someone pleaseee explain these type of problems to me? When is “y” a Function of “x”? 10.Textbook Section 3.2 Exercise # 28 Answer Choices: A. In the given equation, y is a function of x. B. In the given equation, y is not a function of x. 2x-9=6y+2

OpenStudy (anonymous):

lets get the equation in terms of y. \[y = \frac{ 2x-11 }{ 6 }\] Now if y is a function of x then there would only be one value of y for any given value of x... Does this equation work?

OpenStudy (anonymous):

so the answer would be yes?

OpenStudy (anonymous):

because there is only one problem for y?

OpenStudy (anonymous):

if you plug in a number like 1 or 0 for x... does y come out to be just one answer of can y be multiple answers?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

when I solve the problem my answer is -9/6 but I can break that down to 3/2 does that make it two answers?

OpenStudy (anonymous):

no its still one answer because that is 3/2 is just a simplified form of 9/6... which makes it one answer

OpenStudy (anonymous):

so your answer is A... ill brb

OpenStudy (anonymous):

ohhh okay - and I treat every problem like this? For example: how would I solve this one?

OpenStudy (anonymous):

2y2 − 3x2 = 8x

OpenStudy (anonymous):

its 2y^2 and 3x^2

OpenStudy (anonymous):

\[y = \sqrt{\frac{ 8x+3x ^{2} }{ 2 }}\]

OpenStudy (anonymous):

if you look at the graph of this equation it looks like an absolute value function and has only one x value for each y value.